# FedDyMem: Efficient Federated Learning with Dynamic Memory and Memory-Reduce for Unsupervised Image Anomaly Detection

### v1

@@ Abstract
Unsupervised image anomaly detection (UAD) has become a critical process in industrial and medical applications, but it faces growing challenges due to increasing concerns over data privacy. The limited class diversity inherent to one-class classification tasks, combined with distribution biases caused by variations in products across and within clients, poses significant challenges for preserving data privacy with federated UAD. Thus, this article proposes an efficient federated learning method with dynamic memory and memory-reduce for unsupervised image anomaly detection, called FedDyMem. Considering all client data belongs to a single class (i.e., normal sample) in UAD and the distribution of intra-class features demonstrates significant skewness, FedDyMem facilitates knowledge sharing between the client and server through the client’s dynamic memory bank instead of model parameters. In the local clients, a memory generator and a metric loss are employed to improve the consistency of the feature distribution for normal samples, leveraging the local model to update the memory bank dynamically. For efficient communication, a memory-reduce method based on weighted averages is proposed to significantly decrease the scale of memory banks. On the server, global memory is constructed and distributed to individual clients through k-means aggregation. Experiments conducted on six industrial and medical datasets, comprising a mixture of six products or health screening types derived from eleven public datasets, demonstrate the effectiveness of FedDyMem.
Index Terms:
@@
Federated learning, unsupervised anomaly detection, feature distribution shift, communication efficiency, dynamic memory bank.
I Introduction
@@
Unsupervised image anomaly detection has achieved significant success in various domains, such as industrial inspection [1] and medical disease recognition [2]. However, these methods highly depend on the availability of large-scale datasets for centralized training. Industrial companies and healthcare organizations often have practical limitations in collecting and aggregating raw data, which significantly challenges centralized learning approaches [3, 4]. Recently, federated learning has been developed as a privacy-preserving paradigm for machine learning, providing collaborative model training across distributed devices while maintaining raw data locally and not transmitting to central servers [5, 6]. Thus, the integration of local training and global aggregation strategies has garnered significant attention in industrial and healthcare domains, enabling data privacy protection [7, 8, 9].
@@
Federated learning is applicable to both supervised and unsupervised learning scenarios. Recent developments in unsupervised federated learning have shown significant progress[10]. These methods are often focused on learning a general representation or prototype by self-supervised learning (SSL) while keeping private data decentralized and unlabeled. Some methods continue to rely on local updates and aggregation of model parameters to train a global model for representation generation[11, 12]. Alternatively, other approaches take advantage of representation sharing, using knowledge distillation (KD) to construct a robust representation space [13, 14, 15]. However, these methods are designed to obtain a more efficient representation by SSL and require local fine-tuning through a supervised linear evaluation protocol [11] following federated learning. Consequently, existing approaches in unsupervised federated learning are inadequate for addressing the UAD task.
@@
With the advancements in UAD using deep learning, reconstruction-based methods [16, 17, 18] have emerged as prominent strategies by learning to accurately reconstruct normal samples while minimizing reconstruction loss, thereby enabling the identification of anomalies as deviations from expected reconstructions. Intuitively, integrating reconstruction loss functions to locally optimize the model, followed by aggregating the model weights through federated learning approaches, presents a promising avenue for enabling federated learning in UAD. However, as shown in Fig. 1, the intra-class distribution bias within the normal sample class, caused by variations in products, is also an important issue in UAD [19]. Similarly, data heterogeneity in federated learning can lead to significant performance degradation and convergence challenges, primarily due to inconsistencies in local updates and optimization dynamics [20, 21]. Recently, most of the current federated learning methods focus on addressing the challenges posed by data heterogeneity[22, 23, 24, 25, 26, 27]. To address the feature shift challenge in federated learning, some methods have proposed incorporating local constraints by the global model or gradient information, thereby mitigating discrepancies across clients [28, 22, 23, 20, 29]. Despite these efforts, the significant intra-class feature distribution bias in UAD results in insufficient global information to effectively capture the representations and knowledge of diverse products [30]. Meanwhile, other researchers have addressed this issue by synthesizing data [31, 32]. However, sharing additional data among clients inherently increases the risk of data privacy leakage [33]. Recent approaches have increasingly emphasized prototype-learning methods to tackle the challenges posed by inconsistent data distributions. These methods construct category-specific global and local prototypes by leveraging feature alignment, effectively addressing generalization gaps across local clients [34, 35, 27, 26]. Nevertheless, reconstruction-based loss functions in UAD often lead to overfitting on local features of normal samples [36]. The use of a single category hinders effective global feature aggregation and prevents normal representations from achieving a compact distribution. As a result, the aggregated prototypes deviate from the desired global distribution of normal samples.
@@
Furthermore, communication between clients in federated learning is limited by network bandwidth, unreliable connections, and varying device statuses. Consequently, designing communication-efficient algorithms is crucial to overcome these challenges [37]. Existing methods primarily tackle this issue by either decreasing the total number of communication rounds required for model convergence or reducing the data uploaded during each communication round[38]. To accelerate global convergence, some approaches leverage mechanisms such as localized model updates[39] and gradient optimization enhancements[40], effectively reducing the number of communication rounds. To reduce the data uploaded, methods such as model compression [41, 42, 43] and selective client participation [44] have been extensively explored to minimize communication overhead. In this article, we aim to enhance the efficiency of federated learning by minimizing communication overhead through effective data reduction.
@@
To summarize, federated UAD must address the following key challenges: 1) enabling each client to perform end-to-end unsupervised anomaly detection without the reliance on additional labeled data or supervised fine-tuning, utilizing only one-class (normal) samples as the basis for training. 2) mitigating significant distribution biases in normal sample data across clients, which arise due to the heterogeneous product types or health screening modalities within the federated learning framework. 3) ensuring efficient communication between clients and the central server to minimize overhead. This article introduces FedDyMem, an efficient federated learning framework with dynamic memory and memory-reduce, specifically designed to tackle the aforementioned challenges in federated UAD. FedDyMem facilitates knowledge transfer in federated learning by sharing the memory bank instead of model parameters among clients. Specifically, it constructs a local memory bank using the limited dataset of normal samples available at each client. To address the challenge of intra-class distribution bias in different clients, a memory generator and a metric loss function are introduced to improve the consistency of normal feature distributions. Then the local memory bank is dynamically updated by the local client model before uploading. To mitigate distribution bias during aggregation, FedDyMem employs k-means clustering to extract general distribution from different clients, thereby ensuring effective aggregation without confusion. Considering communication efficiency, a memory-reduce method is proposed to decrease the scale of memory banks during dynamic updates. To construct diverse data for our experiments, we categorize the 11 public datasets into six distinct types based on their inherent characteristics. We then evaluate our proposed method separately on six product or health screening types, achieving state-of-the-art performance in each case. The major contributions of this article are summarized as follows:
@@
  • To the best of our knowledge, FedDyMem is the first federated learning framework designed specifically for unsupervised image anomaly detection. This framework facilitates collaborative training using only normal samples available on the client side. To ensure privacy-preserving knowledge sharing, we introduce a dynamic memory bank as an alternative to exchanging model parameters.
@@
  • To improve communication efficiency, we propose a memory-reduce method based on a dynamic weighted average. This method substantially decreases the size of the memory bank, thereby minimizing communication overhead while preserving performance.
@@
II Related Work
II-A Unsupervised Image Anomaly Detection
@@
The objective of UAD is to identify whether a given sample is anomalous and to precisely localize the anomaly regions, using a training dataset that comprises only normal samples[1]. The reconstruction-based methods are significant in UAD, which assumes that anomalous samples cannot be accurately reconstructed by feature learning models trained exclusively on normal samples [45, 46, 47, 17, 48, 18]. In contrast to traditional UAD approaches, which rely on a single, centralized training dataset, federated UAD requires the efficient aggregation of information from samples distributed across multiple clients. Reconstruction-based methods, however, typically rely solely on normal samples within the training set and fail to capture key characteristics of out-of-distribution samples. Another widely utilized approach for anomaly detection is the memory bank-based method. SPADE [49] proposed a semantic pyramid structure to construct a pixel-level feature memory, effectively facilitating anomaly detection. PaDiM [50] employed Gaussian distributions derived from normal samples as the memory bank and utilized the Mahalanobis distance as the anomaly metric. PatchCore [51] introduced local neighborhood aggregation to expand the receptive field of the memory bank while preserving resolution. Additionally, PatchCore implemented a greedy core-set subsampling strategy to reduce the memory bank size without significant performance degradation. PNI [52] further advanced the paradigm by integrating spatial and neighborhood information into the memory bank construction, thereby improving its capacity to represent normal samples comprehensively. These methods extract features from normal images and store them within a feature memory bank. During testing, the sample queries the memory bank to retrieve feature corresponding to the k-nearest neighbors. However, the memory banks are non-trainable, limiting their capacity to learn information across local clients in the federated learning.
II-B Federated Learning
@@
The federated learning can be divided into supervised-based and unsupervised-based methods. In unsupervised federated learning, most existing methods focused on representation learning or prototype-based learning, which aimed to learn good feature representations from unlabeled data to facilitate downstream machine learning tasks [11, 12, 13, 14, 15]. However, these methods depend on aligning local representations with global representations, a task that proves challenging for UADs constrained to a single category. ProtoFL[53] was proposed to integrate federated learning with prototyping for addressing one-class classification. However, its two-stage fine-tuning method, which utilizes normalizing flow for representation learning, introduces unnecessary computational overhead and ignores distribution bias among clients in UAD. Furthermore, most unsupervised federated learning approaches require supervised fine-tuning, which limits their applicability to UAD.
@@
FedAvg [54] is a foundational approach in federated learning that enables the training of a global model by aggregating parameters from locally trained models. Despite its effectiveness, the performance of FedAvg degrades significantly with high data heterogeneity. FedProx [22] and SCAFFOLD [20] demonstrated enhanced performance through a global penalty term, effectively addressing and mitigating discrepancies. Other methods have now been developed to address the data heterogeneity with the personalized model. FedBN [23] mitigates data heterogeneity challenges in federated learning by maintaining BN parameters that are specific to each client’s local model. APPLE[29] aggregates client models locally by learning precise weight updates instead of relying on approximations. Recent approaches address generalization gaps across local clients by leveraging feature alignment to construct category-specific global and local prototypes[34, 27, 26, 35]. Nevertheless, the aforementioned methods face significant challenges in UAD, including high communication overhead, insufficient representation diversity and pronounced intra-class distribution bias.
III Preliminary
III-A Typical Federated Learning
@@
In the typical federated learning, such as FedAvg [54], there exist $N$ clients $\{{\cal C}_{n}\}_{n=1}^{N}$ with their individual datasets ${\{{\cal D}_{n}=({\cal X}_{n},{\cal Y}_{n})\sim{\mathbb{P}}_{n}({\cal X},{\cal Y})\}}_{n=1}^{N}$ and local models $\{f^{n}(\cdot;{\bf W}^{n})\colon{{\cal X}_{n}\to{{\cal Y}_{n}}}\}_{n=1}^{N}$, where $n$ is the index for the client, $({\cal X}_{n},{\cal Y}_{n})$ denotes a set of samples and labels on the $n$-th client, ${{\mathbb{P}_{n}}}({\cal X},{\cal Y})$ represent the joint distribution of samples and labels. The objective of the typical federated learning is to facilitate collaborative training of a global model ${\bf W}^{G}$ that generalizes effectively across the entire dataset $\cal D$ while ensuring data privacy. The global training optimization target is:
$\underset{{\bf W}^{G}}{\min{\cal L}({\bf W}^{G})}\buildrel\Delta\over{=}\sum\limits_{n=1}^{N}{z_{n}{{\cal L}_{n}}({\bf W}^{n})},$ (1)
@@
III-B Federated UAD
As shown in Fig. 1, each client in UDA could be an industrial manufacturing enterprise, an organizational entity or a healthcare institution, etc. that faces the challenges of large raw abnormal data. Consequently, the training dataset available at the $n$-th client is denoted as:

### v2

@@ Abstract
Unsupervised image anomaly detection (UAD) has become a critical process in industrial and medical applications, but it faces growing challenges due to increasing concerns over data privacy. The limited class diversity inherent to one-class classification tasks, combined with distribution biases caused by variations in products across and within clients, poses significant challenges for preserving data privacy with federated UAD. Thus, this article proposes an efficient federated learning method with dynamic memory and memory-reduce for unsupervised image anomaly detection, called FedDyMem. Considering all client data belongs to a single class (i.e., normal sample) in UAD and the distribution of intra-class features demonstrates significant skewness, FedDyMem facilitates knowledge sharing between the client and server through the client’s dynamic memory bank instead of model parameters. In the local clients, a memory generator and a metric loss are employed to improve the consistency of the feature distribution for normal samples, leveraging the local model to update the memory bank dynamically. For efficient communication and data privacy, a memory-reduce method based on weighted averages is proposed to significantly decrease the scale of memory banks. This reduced representation inherently, thereby mitigating the risk of data reconstruction. On the server, global memory is constructed and distributed to individual clients through k-means aggregation. Experiments conducted on six industrial and medical datasets, comprising a mixture of six products or health screening types derived from eleven public datasets, demonstrate the effectiveness of FedDyMem.
keywords:
@@
Unsupervised federated learning, Unsupervised image anomaly detection, Feature distribution shift, Communication efficiency, Privacy-preserving
1 Introduction
@@
Unsupervised image anomaly detection has achieved significant success in various domains, such as industrial inspection ([xie2024iad]) and medical disease recognition ([cai2024medianomaly]). However, these methods highly depend on the availability of large-scale datasets for centralized training. Industrial companies and healthcare organizations often have practical limitations in collecting and aggregating raw data, which significantly challenges centralized learning approaches ([yang2024clustering, wu2024facmic]). Recently, federated learning has been developed as a privacy-preserving paradigm for machine learning, providing collaborative model training across distributed devices while maintaining raw data locally and not transmitting to central servers ([li2021fedrs, saha2024multifaceted]). Thus, the integration of local training and global aggregation strategies has garnered significant attention in industrial and healthcare domains, enabling data privacy protection ([zhao2024medical, zhang2024vertical]).
Federated learning is applicable to both supervised and unsupervised learning scenarios. Recent developments in unsupervised and semi-supervised federated learning have shown significant progress ([jin2023federated, yang2025relation]). These methods are often focused on learning a general representation or prototype by self-supervised learning (SSL) while keeping private data decentralized and unlabeled. Some methods continue to rely on local updates and aggregation of model parameters to train a global model for representation generation ([fedU, liao2024rethinking]). Alternatively, other approaches take advantage of representation sharing, using knowledge distillation (KD) to construct a robust representation space ([fedX, lubana2022orchestra, zhang2024prototype]). However, these methods are designed to obtain a more efficient representation by SSL and require local fine-tuning through a supervised linear evaluation protocol ([fedU]) following federated learning. Consequently, existing approaches in unsupervised federated learning are inadequate for addressing the UAD task.
@@
With the advancements in UAD using deep learning, reconstruction-based methods ([liang2023omni, dai2024generating] have emerged as prominent strategies by learning to accurately reconstruct normal samples while minimizing reconstruction loss, thereby enabling the identification of anomalies as deviations from expected reconstructions. Intuitively, integrating reconstruction loss functions to locally optimize the model, followed by aggregating the model weights through federated learning approaches, presents a promising avenue for enabling federated learning in UAD. However, as shown in Fig. 1, the intra-class distribution bias within the normal sample class, caused by variations in products, is also an important issue in UAD ([lei2025adapted]). Similarly, data heterogeneity in federated learning can lead to significant performance degradation and convergence challenges, primarily due to inconsistencies in local updates and optimization dynamics ([karimireddy2020scaffold, zhao2018federated]). Recently, most of the current federated learning methods focus on addressing the challenges posed by data heterogeneity ([fedprox, li2021fedbn, li2022federated, zhou2024federated, dai2023tackling]). To address the feature shift challenge in federated learning, some methods have proposed incorporating local constraints by the global model or gradient information, thereby mitigating discrepancies across clients ([fedprox, li2021fedbn, karimireddy2020scaffold, luo2022adapt]). Despite these efforts, the significant intra-class feature distribution bias in UAD results in insufficient global information to effectively capture the representations and knowledge of diverse products ([huang2023rethinking]). Meanwhile, other researchers have addressed this issue by synthesizing data ([nguyen2021federated, li2024feature]). However, sharing additional data among clients inherently increases the risk of data privacy leakage ([li2024fedcir]). Recent approaches have increasingly emphasized prototype-learning methods to tackle the challenges posed by inconsistent data distributions. These methods construct category-specific global and local prototypes by leveraging feature alignment, effectively addressing generalization gaps across local clients ([tan2022fedproto, feddbe, dai2023tackling, zhou2024federated]). Nevertheless, reconstruction-based loss functions in UAD often lead to overfitting on local features of normal samples ([you2022unified]). The use of a single category hinders effective global feature aggregation and prevents normal representations from achieving a compact distribution. As a result, the aggregated prototypes deviate from the desired global distribution of normal samples.
@@
Furthermore, communication between clients in federated learning is limited by network bandwidth, unreliable connections, and varying device statuses. Consequently, designing communication-efficient algorithms is crucial to overcome these challenges ([almanifi2023communication]). Existing methods primarily tackle this issue by either decreasing the total number of communication rounds required for model convergence or reducing the data uploaded during each communication round ([li2020federated]). To accelerate global convergence, some approaches leverage mechanisms such as localized model updates ([li2020federated2]) and gradient optimization enhancements ([yang2022federated]), effectively reducing the number of communication rounds. To reduce the data uploaded, methods such as model compression ([shlezinger2020uveqfed, li2023online]) and selective client participation ([xu2020client]) have been extensively explored to minimize communication overhead. Beyond efficiency, privacy-ensured communication is equally essential, especially in resource-constrained and privacy-sensitive environments. The exchanged information should resist reconstruction attempts from adversaries, preventing leakage of the raw client data ([cao2025hybrid]). In this article, we aim to minimize simultaneously communication overhead and the mutual information with the original data through effective data reduction techniques, thereby improving the communication efficiency and enhancing privacy protection.
@@
To summarize, federated UAD must address the following key challenges: 1) enabling each client to perform end-to-end unsupervised anomaly detection without the reliance on additional labeled data or supervised fine-tuning, utilizing only one-class (normal) samples as the basis for training. 2) mitigating significant distribution biases in normal sample data across clients, which arise due to the heterogeneous product types or health screening modalities within the federated learning framework. 3) ensuring communication between clients and the central server is both privacy-preserving and communication-efficient. This article introduces FedDyMem, an efficient federated learning framework with dynamic memory and memory-reduce, specifically designed to tackle the aforementioned challenges in federated UAD. FedDyMem facilitates knowledge transfer in federated learning by sharing the memory bank instead of model parameters among clients. Specifically, it constructs a local memory bank using the limited dataset of normal samples available at each client. To address the challenge of intra-class distribution bias in different clients, a memory generator and a metric loss function are introduced to improve the consistency of normal feature distributions. Then the local memory bank is dynamically updated by the local client model before uploading. To mitigate distribution bias during aggregation, FedDyMem employs k-means clustering to extract general distribution from different clients, thereby ensuring effective aggregation without confusion. Considering communication efficiency and data privacy, a memory-reduce method is proposed to decrease the scale of memory banks, thereby reducing its mutual information with the original data during dynamic updates. To construct diverse data for our experiments, we categorize the 11 public datasets into six distinct types based on their inherent characteristics. We then evaluate our proposed method separately on six product or health screening types, achieving state-of-the-art performance in each case. The major contributions of this article are summarized as follows:
@@
  • To the best of our knowledge, FedDyMem is the first federated learning framework designed specifically for unsupervised image anomaly detection. This framework facilitates collaborative training using only normal samples available on the client side.
@@
  • To achieve both communication efficiency and privacy-preserving knowledge sharing, we propose a memory-reduce method based on a dynamic weighted average. This method substantially decreases the size of the memory bank, thereby reducing both communication overhead and the mutual information between uploaded information and raw data, while preserving model performance.
@@
2 Related Work
2.1 Unsupervised Image Anomaly Detection
@@
The objective of UAD is to identify whether a given sample is anomalous and to precisely localize the anomaly regions, using a training dataset that comprises only normal samples ([xie2024iad]). The reconstruction-based methods are significant in UAD, which assumes that anomalous samples cannot be accurately reconstructed by feature learning models trained exclusively on normal samples ([bergmann2018improving, zavrtanik2021draem, jiang2022masked, liang2023omni, dai2024generating]). In contrast to traditional UAD approaches, which rely on a single, centralized training dataset, federated UAD requires the efficient aggregation of information from samples distributed across multiple clients. Reconstruction-based methods, however, typically rely solely on normal samples within the training set and fail to capture key characteristics of out-of-distribution samples. Another widely utilized approach for anomaly detection is the memory bank-based method. SPADE ([cohen2020sub]) proposed a semantic pyramid structure to construct a pixel-level feature memory, effectively facilitating anomaly detection. PaDiM ([defard2021padim]) employed Gaussian distributions derived from normal samples as the memory bank and utilized the Mahalanobis distance as the anomaly metric. PatchCore ([roth2022towards]) introduced local neighborhood aggregation to expand the receptive field of the memory bank while preserving resolution. Additionally, PatchCore implemented a greedy core-set subsampling strategy to reduce the memory bank size without significant performance degradation. PNI ([bae2023pni]) further advanced the paradigm by integrating spatial and neighborhood information into the memory bank construction, thereby improving its capacity to represent normal samples comprehensively. These methods extract features from normal images and store them within a feature memory bank. During testing, the sample queries the memory bank to retrieve feature corresponding to the k-nearest neighbors. However, the memory banks are non-trainable, limiting their capacity to learn information across local clients in the federated learning.
2.2 Federated Learning
@@
The federated learning can be divided into supervised-based and unsupervised-based methods. In unsupervised federated learning, most existing methods focused on representation learning or prototype-based learning, which aimed to learn good feature representations from unlabeled data to facilitate downstream machine learning tasks ([fedU, liao2024rethinking, fedX, lubana2022orchestra, zhang2024prototype]). However, these methods depend on aligning local representations with global representations, a task that proves challenging for UADs constrained to a single category. ProtoFL ([kim2023protofl]) was proposed to integrate federated learning with prototyping for addressing one-class classification. However, its two-stage fine-tuning method, which utilizes normalizing flow for representation learning, introduces unnecessary computational overhead and ignores distribution bias among clients in UAD. Furthermore, most unsupervised federated learning approaches require supervised fine-tuning, which limits their applicability to UAD.
@@
FedAvg ([fedavg]) is a foundational approach in federated learning that enables the training of a global model by aggregating parameters from locally trained models. Despite its effectiveness, the performance of FedAvg degrades significantly with high data heterogeneity. FedProx ([fedprox]) and SCAFFOLD ([karimireddy2020scaffold]) demonstrated enhanced performance through a global penalty term, effectively addressing and mitigating discrepancies. Other methods have now been developed to address the data heterogeneity with the personalized model. FedBN ([li2021fedbn]) mitigates data heterogeneity challenges in federated learning by maintaining BN parameters that are specific to each client’s local model. APPLE ([luo2022adapt]) aggregates client models locally by learning precise weight updates instead of relying on approximations. Recent approaches address generalization gaps across local clients by leveraging feature alignment to construct category-specific global and local prototypes ([tan2022fedproto, dai2023tackling, zhou2024federated, feddbe]). Nevertheless, the aforementioned methods face significant challenges in UAD, including high communication overhead, insufficient representation diversity and pronounced intra-class distribution bias.
Figure 2: Overview of our proposed FedDyMem framework.
3 Preliminary
3.1 Typical Federated Learning
@@
In the typical federated learning, such as FedAvg ([fedavg]), there exist $N$ clients $\{{\cal C}_{n}\}_{n=1}^{N}$ with their individual datasets ${\{{\cal D}_{n}=({\cal X}_{n},{\cal Y}_{n})\sim{\mathbb{P}}_{n}({\cal X},{\cal Y})\}}_{n=1}^{N}$ and local models $\{f^{n}(\cdot;{\bf W}^{n})\colon{{\cal X}_{n}\to{{\cal Y}_{n}}}\}_{n=1}^{N}$, where $n$ is the index for the client, $({\cal X}_{n},{\cal Y}_{n})$ denotes a set of samples and labels on the $n$-th client, ${{\mathbb{P}_{n}}}({\cal X},{\cal Y})$ represent the joint distribution of samples and labels. The objective of the typical federated learning is to facilitate collaborative training of a global model ${\bf W}^{G}$ that generalizes effectively across the entire dataset $\cal D$ while ensuring data privacy. The global training optimization target is:
$\underset{{\bf W}^{G}}{\min{\cal L}({\bf W}^{G})}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\sum\limits_{n=1}^{N}{z_{n}{{\cal L}_{n}}({\bf W}^{n})},$ (1)
@@
3.2 Federated UAD
As shown in Fig. 1, each client in UAD could be an industrial manufacturing enterprise, an organizational entity or a healthcare institution, etc. that faces the challenges of large raw abnormal data. Consequently, the training dataset available at the $n$-th client is denoted as: